OpenStudy (anonymous):

please. help. me. m^-5 = 1/32 m=?

OpenStudy (anonymous):

@DarkMoonZ @Anon101 @fierstar123 @HelpBlahBlahBlah @Callisto @xo.minnie.xox @Michele_Laino @thomaster

OpenStudy (anonymous):

finally o.o

OpenStudy (michele_laino):

hint: we rewrite your equation as follows: \[\Large {m^{ - 5}} = \frac{1}{{{m^5}}} = {\left( {\frac{1}{m}} \right)^5} = \frac{1}{{32}}\]

OpenStudy (anonymous):

so the answer is 32 XD

OpenStudy (michele_laino):

are you sure?

OpenStudy (anonymous):

umm no D:

OpenStudy (michele_laino):

hint: \[\Large \frac{1}{{32}} = \frac{1}{{{2^5}}} = {\left( {\frac{1}{2}} \right)^5}\]

OpenStudy (michele_laino):

so your equation, becomes: \[\Large {\left( {\frac{1}{m}} \right)^5} = {\left( {\frac{1}{2}} \right)^5}\]

OpenStudy (anonymous):

ohhhh okay.

OpenStudy (anonymous):

Seems legit.

OpenStudy (anonymous):

but what else after that

OpenStudy (anonymous):

bc u cant simplify 1/2

OpenStudy (michele_laino):

two powers with the same base, are equal each other when the bases are equal each other

OpenStudy (michele_laino):

so we can write: \[\Large \frac{1}{m} = \frac{1}{2}\] what is m?

OpenStudy (michele_laino):

oops... two powers with the same exponent, are equal each other when the bases are equal each other

OpenStudy (anonymous):

sorry my computer is going crazy

OpenStudy (michele_laino):

so, what is m?

OpenStudy (anonymous):

2

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

THANKYOUUUU XD

OpenStudy (michele_laino):

thank you! :)

OpenStudy (anonymous):

ily XD <3